OpenStudy (lizilizi28):

Radical Expressions 3√64(x-2)^3 simplify the expression assumme that all variables are real number

OpenStudy (anonymous):

is 24(x-2)^3 a option?

jimthompson5910 (jim_thompson5910):

Similar Problem \[\Large \sqrt[3]{8(x-5)^3}\] \[\Large \sqrt[3]{8}\sqrt[3]{(x-5)^3}\] \[\Large \sqrt[3]{2^3}\sqrt[3]{(x-5)^3}\] \[\Large 2(x-5)\] \[\Large 2x-10\] So \[\Large \sqrt[3]{8(x-5)^3}\] simplifies to \[\Large 2x-10\] Note: this is NOT your answer as it's just an example

OpenStudy (lizilizi28):

i think my aswer can be 2x-4 im correct

OpenStudy (anonymous):